OpenStudy (anonymous):

The voltage V (volts), current I (amperes), and resistence R (ohms) of an electric circuit like the one shown on p. 192, #15 are related by the equation V = IR. Suppose that V is increasing at a rate of 1 volt/sec while I is decreasing at a rate of 1/7 amp/sec. Let t denote time in seconds. Determine the rate at which R is changing when V = 1 volts and I = 5 amperes.

OpenStudy (perl):

Does it say to use calculus?

OpenStudy (anonymous):

Yeah this is calc

OpenStudy (anonymous):

All I need help with is the differentiating the equation, do I use the multiplication rule? Or no

OpenStudy (perl):

yes

OpenStudy (perl):

product rule

OpenStudy (anonymous):

I have been working through the problem, but I have not complete calculus yet. Good luck.

OpenStudy (perl):

$$ \Large \frac {dV}{dt} = \frac{dI}{dt}R+ I\frac{dR}{dt} $$

OpenStudy (anonymous):

Awesome thanks! Wasn't sure if it was different for a problem like this